OpenStudy (marziajackson$):

Javi figures that 2 small pizzas will feed 3 people. He expects 15 people at the team pizza party. How many small pizzas will he need for 15 people? A. 6 B. 10 C. 30 D. 45

OpenStudy (marziajackson$):

@Kstate Can you help me?

OpenStudy (anonymous):

Hey hun whats up

OpenStudy (marziajackson$):

Good

OpenStudy (anonymous):

A

OpenStudy (marziajackson$):

Okay Thank you

OpenStudy (anonymous):

was i right

OpenStudy (marziajackson$):

yes

OpenStudy (anonymous):

pellet no its 10 sorry

OpenStudy (texaschic101):

2/3 = x/15 cross multiply 3x = 30 x = 10

OpenStudy (anonymous):

i was right

OpenStudy (anonymous):

yah but @texaschic101 if 2 pizzas feed 3 people u would only need 6 because if u draw it out which i will do for u u will see

OpenStudy (anonymous):

theres how u would get the answer by the way